the difference between 2 numbers is three if the square of the smaller number is equal to 4 times the larger number what is the number?
Found 2 solutions by stanbon, Cromlix:
Answer by stanbon(75887)

the difference between 2 numbers is three if the square of the smaller number is equal to 4 times the larger number what is the number?
------
Equations:
x - y = 3
y^2 = 4x
------
Substitute for "x" and solve for "y"::
(y^2/4) - y = 3
-----
y^2 -4y - 12 = 0
Factor::
(y-6)(y+2) = 0
If y = 6, x = 9
If y = -2, x = 1
--------------
Cheers,
Stan H.
------------

Answer by Cromlix(4381)

Hi there,
the difference between 2 numbers is three
x - y = 3
x = 3 + y
the square of the smaller number is equal
to 4 times the larger number.
y^2 = 4x
Substitute x with (3 + y) from top equation
y^2 = 4(3 + y)
y^2 = 12 + 4y
Rearrange into ay^2 + by + c form
y^2 - 4y - 12 = 0
(y + 2)(y - 6)= 0
y + 2 = 0
y = -2 (Not used as -ve)
y - 6 = 0
y = 6
Substitute y = 6 into:
x - y = 3
x - 6 = 3
x = 3 + 6
x = 9.
Smaller number = 6
Larger number = 9
Hope this helps :-)